"5 times the sum of 3 consecutive integers equal 150. what are the integers?
Answer by algebrahouse.com(1659)

"5 times the sum of 3 consecutive integers equal 150. what are the integers?"

x = first integer
x + 1 = second integer
x + 2 = third integer

5(x + x + 1 + x + 2) = 150 {five times sum is 150}
5(3x + 3) = 150 {combined like terms in parentheses}
15x + 15 = 150 {used distributive property}
15x = 135 {subtracted 15 from both sides}
x = 9 {divided both sides by 15}
x + 1 = 10 {substituted 9, in for x, into x + 1}
x + 2 = 11 {substituted 9, in for x, into x + 2}

9, 10, and 11 are the three consecutive integers
www.algebrahouse.com